# Physics beyond the Standard Model with the DSA-2000

### v1

@@ Abstract
The upcoming Deep Synoptic Array 2000 (DSA-2000) will map the radio sky at $0.7-2$ GHz ($2.9-8.3\,\mu$eV) with unprecedented sensitivity. This will enable searches for dark matter and other physics beyond the Standard Model, of which we study four cases: axions, dark photons, dark matter subhalos and neutrino masses. We forecast DSA-2000’s potential to detect axions through two mechanisms in neutron star magnetospheres: photon conversion of axion dark matter and radio emission from axion clouds, developing the first analytical treatment of the latter. We also forecast DSA-2000’s sensitivity to discover kinetically mixed dark photons from black hole superradiance, constrain dark matter substructure and fifth forces through pulsar timing, and improve cosmological neutrino mass inference through fast radio burst dispersion measurements. Our analysis indicates that in its planned five year run the DSA-2000 could reach sensitivity to QCD axion parameters, improve current limits on compact dark matter by an order of magnitude, and enhance cosmological weak lensing neutrino mass constraints by a factor of three.
@@ 1 Introduction
The hunt for physics beyond the Standard Model (BSM) requires diverse observational approaches spanning different scales and wavelengths. Despite extensive searches, fundamental questions remain about the nature of dark matter (DM), neutrino masses, and potential new forces and particles. Radio astronomy offers unique windows into these questions, with the upcoming Deep Synoptic Array 2000 (DSA-2000) representing a considerable advancement in our observational capabilities [1].
The DSA-2000 will be an interferometric array of 2000 × 5 m antennas to be built in a radio-quiet valley in Nevada, USA. Operating at 0.7-2.0 GHz ($2.9-8.3\,\mu$eV), its notable feature is the combination of small dishes that provide a large field of view together with substantial collecting area, resulting in substantial survey speed advantages for specific observing scenarios. For example, the DSA-2000 can image a 10 deg2 field to 1 $\mu$Jy in one hour, while the proposed mid-frequency Square Kilometer Array (SKA1-Mid) would require six hours, and the Karl G. Jansky Very Large Array (VLA) would need approximately 50 days of continuous observing.
@@
We present forecasts for multiple BSM scenarios, distinguishing between searches that can be conducted within the planned DSA-2000 survey program and those requiring additional dedicated observations. The all-sky survey, which is part of the core observing strategy, will enable searches for dark photon superradiance signals from black holes [2], and improve neutrino mass constraints through FRB dispersion measurements [3]. We forecast that DSA-2000 will improve cosmological neutrino mass constraints from weak lensing measurements by a factor of three through breaking degeneracies with baryonic feedback. For axion searches, we consider two distinct scenarios: (1) radio signals from axion DM conversion in neutron star (NS) magnetospheres [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], for which we advocate dedicated 10-hour high-resolution beamforming observations of select magnetars. Beamforming refers to phasing up the array in a particular direction, in this case preserving a single spatial pixel at the position of a known NS, but with very high time and frequency resolution. And (2), radio signals from axion clouds that form around NSs [28, 29], which we can search for using both the planned all-sky survey data and targeted beamforming observations, potentially reaching sensitivity to the QCD axion band. Here, we derive the first analytical description of radio signals from axion clouds around NSs as a function of period and magnetic field—a signal previously quantified only through simulations [28]. This analytical approach captures key features of the full simulation while providing deeper physical insight into the underlying processes, enabling more efficient exploration of the parameter space relevant to DSA-2000 observations.
Lastly, the planned discovery of $\sim 130$ new millisecond pulsars suitable for pulsar timing through the DSA-2000 pulsar survey [1] will improve constraints on DM substructure and fifth forces by approximately an order of magnitude compared to current limits [30]. Our analysis shows that while DSA-2000’s planned 5-year operation will yield significant advances, extending millisecond pulsar timing observations over a 25-year period would dramatically enhance sensitivity to DM substructure, potentially reaching levels corresponding to a significant fraction of the local DM density—a threshold that would make this approach particularly powerful for constraining theories of DM.
@@ 2 Radio signals from axions
Axions are hypothesized particles arising from well-motivated extensions of the Standard Model (SM) and are leading DM candidates. With mass $m_{a}\lesssim 1$ eV and high occupation numbers, axion DM is well-described by a classical wave oscillating at frequency $m_{a}/(2\pi)$. The mass range $m_{a}\lesssim$ meV is particularly compelling, as the observed DM density can be produced through non-thermal mechanisms like vacuum misalignment [31, 32, 33] and decay of topological defects [34, 35, 36, 37, 38, 39, 40, 41]. The quantum chromodynamics (QCD) axion is especially well-motivated, as it not only offers a natural explanation for DM but also elegantly resolves the strong CP problem in QCD [42, 43, 44, 45]. The strong CP problem concerns the seemingly unnatural absence of charge-parity (CP) violation in QCD, as evidenced by the stringent experimental limits on the neutron’s electric dipole moment [46]. Introducing the axion field $a(x)$ with a discrete shift symmetry, coupling to QCD via
@@
with $E/N=8/3$ (DFSZ [47, 48]) or $E/N=0$ (KSVZ [49, 50]) in benchmark models. Beyond QCD axions, general axions arise ubiquitously in UV completions and string theory [51, 52, 53, 54]. Both DM and non-DM axions are theoretically motivated, and we explore DSA-2000’s discovery potential for radio signals from axions in both contexts.
2.1 Axion dark matter conversion in NS magnetospheres
@@
NSs, the collapsed cores of massive stars, possess some of the largest magnetic fields in the observed universe, with surface field strengths as high as $10^{15}$ G. Additionally, rotating NSs—pulsars—are surrounded by a dense plasma magnetosphere, consisting of electrons and positrons co-rotating with the NS [^1] Importantly, co-rotation cannot be sustained everywhere in the magnetosphere. The implications of breakdown of co-rotation are expounded in Sec. 2.4.[55], which can further enhance the mixing between axions and photons. A promising avenue towards indirect detection of axion DM is to search for bright, narrow radio lines associated with axion-to-photon conversion in NS magnetospheres [4, 5, 6, 9, 7, 8, 10, 56, 57, 15, 17, 18, 19].
@@
The expected flux of photons from axion conversion within an NS magnetosphere has been extensively derived in the literature [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 58, 20, 59, 21, 22, 28, 29, 60, 23, 24, 25, 26]. Here we briefly summarize the assumptions and derivations.
The electron plasma distribution in the NS magnetosphere is well-described by the Goldreich-Julian (GJ) model [55]. The GJ model provides the minimum electron/positron number density required to screen any accelerating electric fields, which is derived by solving Maxwell’s equations for particles on the magnetic field lines rotating around the star:
@@
where $e$ is the positron charge, $r$ and $\theta$ are the radial distance (measured from the center of the NS) and polar angle in spherical coordinates, and $\boldsymbol{\Omega}$ (taken to be parallel to $\hat{z}$) and $\mathbf{B}(\mathbf{x})$ are the NS rotation frequency vector and magnetic field, respectively. We use the approximation $(\Omega r)^{2}=4\times 10^{-8}(\Omega/(2\pi\,\mathrm{Hz}))^{2}(r/(10\,\mathrm{km}))^{2}\ll 1$ for the second equality, which is valid for the range of NS rotational frequencies and radial distances of interest. The argument of Goldreich and Julian provides a condition on the net charge density, $\rho=-2\boldsymbol{\Omega}\cdot\mathbf{B}(\mathbf{x})$, required to screen the rotationally-generated electric field, with Eq. (2.3) indicating the minimum number density of charges required to provide this charge density. It may be possible that the same charge density is provided by a much larger number density of particles, $n=\kappa n_{e}$, where $\kappa$ is known as the “multiplicity” of the magnetosphere (which may itself be position dependent). High multiplicity is associated with regions of active pair creation. For example, the region supporting open field lines is expected to host a very high multiplicity, possibly reaching $10^{5}$ [61, 62]. By contrast, the closed magnetosphere is expected to be relatively dormant, which supports the assertion that the pair multiplicity should be relatively low therein. This is supported by 3D global simulations, although determining the exact closed-zone multiplicity is an active area of research. For a dipole magnetic field, the $z$-component is
@@
The surface magnetic field of a generic NS can be computed by equating the radiation power of a rotating dipole with the change in NS rotational kinetic energy, giving [63, 64]
@@
Associated with the GJ density in (2.3) is a plasma frequency, $\omega_{p}(\mathbf{x})=\sqrt{e^{2}n_{e}(\mathbf{x})/m_{e}}$ (where $m_{e}$ is the electron mass), which can be understood as an effective photon mass. When the plasma frequency is comparable to the axion mass, i.e. $m_{a}\sim\omega_{p}$, axions can resonantly convert into photons. In vacuum, axion-photon mixing suffers from a mismatch in the dispersion relations of the massive axion and the massless photon. In a magnetized plasma, the axion mixes with a particle eigenmode, called the O-mode, whose dispersion relation is (see, e.g., [56])
@@
$\approx 36\ {\rm km}\left({B_{0}\over 10^{12}\ {\rm G}}\right)^{1/3}\left({\Omega/2\pi\over{\rm Hz}}\right)^{1/3}\left({m_{a}\over\mu{\rm eV}}\right)^{-2/3}[3\cos\theta\mathbf{\hat{m}}\cdot\mathbf{\hat{r}}-\cos\theta_{m}]^{1/3},$ (2.7)
@@
The radiated power of photons depends on the DM density near $r_{c}$, which is enhanced by the NS’s gravitational field. Assuming that DM velocity satisfies the Maxwell-Boltzmann distribution with virial velocity of $v_{0}$, by conservation of energy, the DM mass density near $r_{c}$ is related to its asymptotic DM density at infinity, $\rho_{\mathrm{DM}}$, by [6]
@@
where $M_{\mathrm{NS}}$ is the NS mass. The asymptotic DM mass density $\rho_{\mathrm{DM}}$ is dependent on its position within the galaxy. In this work, for simplicity, we model the DM density profile with the Navarro-Frenk-White (NFW) profile, given by [65, 66]
@@
where $R_{\mathrm{GC}}$ is the distance from the galactic center, $r_{s}\approx 8$ kpc is the scale radius of the profile [67], and $\rho_{0}$ is a normalization constant set to produce $\rho_{\mathrm{DM}}(R_{\odot})=0.4$ GeV/cm3 at the solar radius, $R_{\odot}=8$ kpc.
During most of the axion’s trajectory through the magnetosphere, the dispersion relation of the axion $\omega^{2}=k^{2}+m_{a}^{2}$ differs from that of the O-mode (2.6), with the momentum mismatch $\delta k\equiv|\mathbf{k}_{a}-\mathbf{k}_{\gamma}|\gg L^{-1}$, where $L$ is a typical length scale associated with the NS. Therefore, axion-photon mixing is highly suppressed. On resonance, when $\delta k\ll L^{-1}$, the conversion probability can be derived in the stationary-phase approximation (see, e.g., [6, 56] for a derivation). In the non-relativistic limit, the conversion probability is [68]
@@
where $v_{c}\approx\sqrt{2GM_{\mathrm{NS}}/r_{c}}$ is the DM velocity at $r_{c}$, and $\theta_{c}$ is the angle between the magnetic field and the photon momentum at the conversion point. For simplification, we compute the conversion probability at the equator, where the angle between radially infalling axions and the magnetic field is $\theta_{c}=\pi/2$, and adopt the GJ model, wherein $\omega_{p}(r)\propto r^{-{3/2}}$. Since the conversion occurs resonantly in a small region around $r_{c}$, the radiated power of converted photons from the NS is obtained by considering a flux of DM through a surface with solid angle $d\Omega_{\mathrm{ang}}$ at $r=r_{c}$, i.e. $d\mathcal{P}\approx 2P_{a\to\gamma}\rho^{r_{c}}_{\mathrm{DM}}v_{c}r_{c}^{2}d\Omega_{\mathrm{ang}}$, and is hence given by (cf. Eqs. (2.4)-(2.8) and Eq. (2.10)) [69]
@@
The time-dependence of the power enters in the $\mathbf{\hat{m}}\cdot\mathbf{\hat{r}}$ term, which is periodic with a period that equals the pulsar period $P$. For simplicity in our forecast, we consider the time-averaged signal. However, we note that since the timing resolution of DSA-2000 exceeds the frequency of a typical NS with a large magnetic field (and $P\gtrsim 1$ s), an improvement in the sensitivity by roughly a factor of two can be expected when weighing the signal across the rotation phase.
@@
$=$ $3.1\times 10^{-4}\,\mathrm{mJy}\left(\frac{g_{a\gamma\gamma}}{10^{-12}\,\mathrm{GeV}^{-1}}\right)^{2}\left(\frac{R_{\mathrm{NS}}}{10\,\mathrm{km}}\right)^{5/2}\left(\frac{m_{a}/2\pi}{\mathrm{GHz}}\right)^{1/3}\left(\frac{B_{0}}{10^{12}\,\mathrm{G}}\right)^{5/6}\left(\frac{P}{\mathrm{sec}}\right)^{7/6}$
$\left(\frac{\rho_{\mathrm{DM}}}{0.4\,\mathrm{GeV\,cm}^{-3}}\right)\left(\frac{M_{\mathrm{NS}}}{M_{\odot}}\right)^{1/2}\left(\frac{200\,\mathrm{km/s}}{v_{0}}\right)^{3}\left(\frac{1\,\mathrm{kpc}}{d}\right)^{2}|3\cos\theta\mathbf{\hat{m}}\cdot\mathbf{\hat{r}}-\cos\theta_{m}|^{5/6}\,,$ (2.12)
where $d$ is the Earth-pulsar distance and
$\mathcal{B}\sim\left(\frac{v_{0}}{c}\right)^{2}\left(\frac{m_{a}}{2\pi}\right)=3.4\,\text{kHz}\left(\frac{v_{0}}{200\,\mathrm{km/s}}\right)^{2}\left(\frac{m_{a}/2\pi}{\mathrm{GHz}}\right)\,,$ (2.13)
is the signal bandwidth [^2] Interactions of the axion-sourced photons with magnetospheric plasma may lead to broadening to a bandwidth of $\sim 10^{-5}m_{a}$ [70, 13, 20], but we neglect this effect here for simplicity.. The signal-to-noise ratio (SNR) is given by
@@
Figure 3: Expected radio spectral flux density from axion-photon conversions in NSs, assuming $m_{a}/(2\pi)=1$ GHz and $g_{a\gamma\gamma}=10^{-14}$ GeV-1, for each pulsar in the Australia Telescope National Facility pulsar catalog [71], magnetars in the McGill Magnetar Catalog [72], the Galactic Center magnetar SGR J1745-2900 with $P=3.76$ s [73, 74, 75], and four recently identified long-period magnetars, including PSR J0901-4046 with $P=75.89$ s [76], GLEAM-X J162759.5-523504.3 with $P=18.2$ min [77], GPM J1839-10 with $P=22.0$ min [78], and ASKAP J1935+2148, with $P=53.8$ min [79]. Pulsars and magnetars in the catalogs are denoted by dots, while long-period magnetars are denoted by stars. The background contours in grey are derived assuming $M_{\mathrm{NS}}=1\,M_{\odot}$, $R_{\mathrm{NS}}=10$ km, $\theta=120^{\circ}$, $\theta_{m}=10^{\circ}$, $\rho_{\mathrm{DM}}=0.4$ GeV/cm3, $d=1$ kpc, and $B_{0}$ satisfying Eq. (2.5). The NSs that are not expected to produce any observable flux with these parameters are colored in black. Here we note that PSR J0901-4046 and GLEAM-X J162759.5-523504.3 are not within DSA-2000’s field of view.
@@
To detect a photon signal from axion-photon conversion, one could either target specific NSs expected to exhibit a strong axion-induced flux or measure the total flux from a population of NSs (see, e.g., Ref. [7]). However, in the case of DSA-2000, achieving the required frequency resolution to detect the DM axion signal (cf. Eq. (2.13)) necessitates beamforming with high temporal and spectral resolution on specific sky positions. This limitation makes a collective observation of an NS population unfavorable. Therefore, for the remainder of this Section, we focus on identifying individual targets that offer the most promising sensitivity to DM axion signals.

### v2

@@ Abstract
[…]
@@ 1 Introduction
The hunt for physics beyond the Standard Model (BSM) requires diverse observational approaches spanning different scales and wavelengths. Despite extensive searches, fundamental questions remain about the nature of dark matter (DM), neutrino masses, and potential new forces and particles. Radio astronomy offers unique windows into these questions, with the upcoming Deep Synoptic Array 2000 (DSA-2000) representing a considerable advancement in our observational capabilities [2019BAAS...51g.255H].
[…]
@@
We present forecasts for multiple BSM scenarios, distinguishing between searches that can be conducted within the planned DSA-2000 survey program and those requiring additional dedicated observations. The all-sky survey, which is part of the core observing strategy, will enable searches for dark photon superradiance signals from black holes [Siemonsen:2022ivj], and improve neutrino mass constraints through FRB dispersion measurements [reischke_cosmological_2023]. We forecast that DSA-2000 will improve cosmological neutrino mass constraints from weak lensing measurements by a factor of three through breaking degeneracies with baryonic feedback. For axion searches, we consider two distinct scenarios: (1) radio signals from axion DM conversion in neutron star (NS) magnetospheres [Pshirkov:2007st, Huang:2018lxq, Hook:2018iia, Safdi:2018oeu, Battye:2019aco, Leroy:2019ghm, Foster:2020pgt, Prabhu:2020yif, Buckley:2020fmh, Witte:2021arp, Battye:2021xvt, battye2021robust, Nurmi:2021xds, Foster:2022fxn, Witte:2022cjj, Battye:2023oac, McDonald:2023shx, Xue:2023ejt, Tjemsland:2023vvc, Witte:2024akb, Kouvaris:2022guf, Maseizik:2024qly, Song:2024rru, Roy:2025mqw], for which we advocate dedicated 10-hour high-resolution beamforming observations of select magnetars. Beamforming refers to phasing up the array in a particular direction, in this case preserving a single spatial pixel at the position of a known NS, but with very high time and frequency resolution. And (2), radio signals from axion clouds that form around NSs [Noordhuis:2023wid, Caputo:2023cpv], which we can search for using both the planned all-sky survey data and targeted beamforming observations, potentially reaching sensitivity to the QCD axion band. Here, we derive the first analytical description of radio signals from axion clouds around NSs as a function of period and magnetic field—a signal previously quantified only through simulations [Noordhuis:2023wid]. This analytical approach captures key features of the full simulation while providing deeper physical insight into the underlying processes, enabling more efficient exploration of the parameter space relevant to DSA-2000 observations.
Lastly, the planned discovery of $\sim 130$ new millisecond pulsars suitable for pulsar timing through the DSA-2000 pulsar survey [2019BAAS...51g.255H] will improve constraints on DM substructure and fifth forces by approximately an order of magnitude compared to current limits [NANOGrav:2023hvm]. Our analysis shows that while DSA-2000’s planned 5-year operation will yield significant advances, extending millisecond pulsar timing observations over a 25-year period would dramatically enhance sensitivity to DM substructure, potentially reaching levels corresponding to a significant fraction of the local DM density—a threshold that would make this approach particularly powerful for constraining theories of DM.
@@ 2 Radio signals from axions
Axions are hypothesized particles arising from well-motivated extensions of the Standard Model (SM) and are leading DM candidates. With mass $m_{a}\lesssim 1$ eV and high occupation numbers, axion DM is well-described by a classical wave oscillating at frequency $m_{a}/(2\pi)$. The mass range $m_{a}\lesssim$ meV is particularly compelling, as the observed DM density can be produced through non-thermal mechanisms like vacuum misalignment [PhysRevLett.40.223, DINE1983137, PRESKILL1983127] and decay of topological defects [Battye:1993jv, PhysRevLett.73.2954, DAVIS1989167, Marsh_2016, Gorghetto:2018myk, Gorghetto:2020qws, Buschmann:2019icd, Buschmann:2021sdq]. The quantum chromodynamics (QCD) axion is especially well-motivated, as it not only offers a natural explanation for DM but also elegantly resolves the strong CP problem in QCD [Peccei:1977hh, Peccei:1977ur, PhysRevLett.40.223, Wilczek:1977pj]. The strong CP problem concerns the seemingly unnatural absence of charge-parity (CP) violation in QCD, as evidenced by the stringent experimental limits on the neutron’s electric dipole moment [Abel:2020pzs]. Introducing the axion field $a(x)$ with a discrete shift symmetry, coupling to QCD via
@@
with $E/N=8/3$ (DFSZ [DINE1981199, Zhitnitsky:1980tq]) or $E/N=0$ (KSVZ [PhysRevLett.43.103, SHIFMAN1980493]) in benchmark models. Beyond QCD axions, general axions arise ubiquitously in UV completions and string theory [Svrcek:2006yi, Green_Schwarz_Witten_2012, PhysRevD.81.123530, Demirtas:2021gsq]. Both DM and non-DM axions are theoretically motivated, and we explore DSA-2000’s discovery potential for radio signals from axions in both contexts.
2.1 Axion dark matter conversion in neutron star magnetospheres
@@
NSs, the collapsed cores of massive stars, possess some of the largest magnetic fields in the observed universe, with surface field strengths as high as $10^{15}$ G. Additionally, rotating NSs—pulsars—are surrounded by a dense plasma magnetosphere, consisting of electrons and positrons co-rotating with the NS [^1] Importantly, co-rotation cannot be sustained everywhere in the magnetosphere. The implications of breakdown of co-rotation are expounded in Sec. 2.4.[1969ApJ...157..869G], which can further enhance the mixing between axions and photons. A promising avenue towards indirect detection of axion DM is to search for bright, narrow radio lines associated with axion-to-photon conversion in NS magnetospheres [Pshirkov:2007st, Huang:2018lxq, Hook:2018iia, Leroy:2019ghm, Safdi:2018oeu, Battye:2019aco, Foster:2020pgt, Witte:2021arp, millar2021axionphotonUPDATED, battye2021robust, Foster:2022fxn, Witte:2022cjj, Battye:2023oac], via the process depicted in Fig. 1.
@@
The expected flux of photons from axion conversion within an NS magnetosphere has been extensively derived in the literature [Pshirkov:2007st, Huang:2018lxq, Hook:2018iia, Safdi:2018oeu, Battye:2019aco, Leroy:2019ghm, Foster:2020pgt, Prabhu:2020yif, Buckley:2020fmh, Witte:2021arp, Battye:2021xvt, battye2021robust, Nurmi:2021xds, Foster:2022fxn, Witte:2022cjj, Battye:2023oac, Prabhu:2023cgb, McDonald:2023shx, noordhuis2023novel, Xue:2023ejt, Tjemsland:2023vvc, Noordhuis:2023wid, Caputo:2023cpv, Khelashvili:2024sup, Witte:2024akb, Kouvaris:2022guf, Maseizik:2024qly, Song:2024rru]. Here we briefly summarize the assumptions and derivations.
The electron plasma distribution in the NS magnetosphere is well-described by the Goldreich-Julian (GJ) model [1969ApJ...157..869G]. The GJ model provides the minimum electron/positron number density required to screen any accelerating electric fields, which is derived by solving Maxwell’s equations for particles on the magnetic field lines rotating around the star:
@@
where $e$ is the positron charge, $r$ and $\theta$ are the radial distance (measured from the center of the NS) and polar angle in spherical coordinates, and $\boldsymbol{\Omega}$ (taken to be parallel to $\hat{z}$) and $\mathbf{B}(\mathbf{x})$ are the NS rotation frequency vector and magnetic field, respectively. We use the approximation $(\Omega r)^{2}=4\times 10^{-8}(\Omega/(2\pi\,\mathrm{Hz}))^{2}(r/(10\,\mathrm{km}))^{2}\ll 1$ for the second equality, which is valid for the range of NS rotational frequencies and radial distances of interest. The argument of Goldreich and Julian provides a condition on the net charge density, $\rho=-2\boldsymbol{\Omega}\cdot\mathbf{B}(\mathbf{x})$, required to screen the rotationally-generated electric field, with Eq. (2.3) indicating the minimum number density of charges required to provide this charge density. It may be possible that the same charge density is provided by a much larger number density of particles, $n=\kappa n_{e}$, where $\kappa$ is known as the “multiplicity” of the magnetosphere (which may itself be position dependent). High multiplicity is associated with regions of active pair creation. For example, the region supporting open field lines is expected to host a very high multiplicity, possibly reaching $10^{5}$ [timokhinharding2015, timokhinharding2018]. By contrast, the closed magnetosphere is expected to be relatively dormant, which supports the assertion that the pair multiplicity should be relatively low therein. This is supported by 3D global simulations, although determining the exact closed-zone multiplicity is an active area of research. For a dipole magnetic field, the $z$-component is
@@
The surface magnetic field of a generic NS can be computed by equating the radiation power of a rotating dipole with the change in NS rotational kinetic energy, giving [2012hpa..book.....L, Philippov:2013aha]
@@
Associated with the GJ density in (2.3) is a plasma frequency, $\omega_{p}(\mathbf{x})=\sqrt{e^{2}n_{e}(\mathbf{x})/m_{e}}$ (where $m_{e}$ is the electron mass), which can be understood as an effective photon mass. When the plasma frequency is comparable to the axion mass, i.e. $m_{a}\sim\omega_{p}$, axions can resonantly convert into photons. In vacuum, axion-photon mixing suffers from a mismatch in the dispersion relations of the massive axion and the massless photon. In a magnetized plasma, the axion mixes with a particle eigenmode, called the O-mode, whose dispersion relation is (see, e.g., [Witte2021])
@@
$\approx 14\ {\rm km}\left(\frac{R_{\mathrm{NS}}}{10\,\mathrm{km}}\right)\left(\frac{B_{0}}{10^{12}\,\mathrm{G}}\right)^{1/3}\left(\frac{\mathrm{sec}}{P}\right)^{1/3}\left(\frac{\mathrm{GHz}}{m_{a}/2\pi}\right)^{2/3}[3\cos\theta\mathbf{\hat{m}}\cdot\mathbf{\hat{r}}-\cos\theta_{m}]^{1/3}\,.$ (2.7)
@@
The radiated power of photons depends on the DM density near $r_{c}$, which is enhanced by the NS’s gravitational field. Assuming that DM velocity satisfies the Maxwell-Boltzmann distribution with virial velocity of $v_{0}$, by conservation of energy, the DM mass density near $r_{c}$ is related to its asymptotic DM density at infinity, $\rho_{\mathrm{DM}}$, by [Hook:2018iia]
@@
where $M_{\mathrm{NS}}$ is the NS mass. The asymptotic DM mass density $\rho_{\mathrm{DM}}$ is dependent on its position within the galaxy. In this work, for simplicity, we model the DM density profile with the Navarro-Frenk-White (NFW) profile, given by [Navarro:1995iw, Navarro:1996gj]
@@
where $R_{\mathrm{GC}}$ is the distance from the galactic center, $r_{s}\approx 8$ kpc is the scale radius of the profile [2019MNRAS.487.5679L], and $\rho_{0}$ is a normalization constant set to produce $\rho_{\mathrm{DM}}(R_{\odot})=0.4$ GeV/cm3 at the solar radius, $R_{\odot}=8$ kpc.
During most of the axion’s trajectory through the magnetosphere, the dispersion relation of the axion $\omega^{2}=k^{2}+m_{a}^{2}$ differs from that of the O-mode (2.6), with the momentum mismatch $\delta k\equiv|\mathbf{k}_{a}-\mathbf{k}_{\gamma}|\gg L^{-1}$, where $L$ is a typical length scale associated with the NS. Therefore, axion-photon mixing is highly suppressed. On resonance, when $\delta k\ll L^{-1}$, the conversion probability can be derived in the stationary-phase approximation (see, e.g., [Hook:2018iia, Witte2021] for a derivation). In the non-relativistic limit, the conversion probability is [McDonald:2023ohd]
@@
where $v_{c}\approx\sqrt{2GM_{\mathrm{NS}}/r_{c}}$ is the DM velocity at $r_{c}$, and $\theta_{c}$ is the angle between the magnetic field and the photon momentum at the conversion point. For simplification, we compute the conversion probability at the equator, where the angle between radially infalling axions and the magnetic field is $\theta_{c}=\pi/2$, and adopt the GJ model, wherein $\omega_{p}(r)\propto r^{-{3/2}}$. Since the conversion occurs resonantly in a small region around $r_{c}$, the radiated power of converted photons from the NS is obtained by considering a flux of DM through a surface with solid angle $d\Omega_{\mathrm{ang}}$ at $r=r_{c}$, i.e. $d\mathcal{P}\approx 2P_{a\to\gamma}\rho^{r_{c}}_{\mathrm{DM}}v_{c}r_{c}^{2}d\Omega_{\mathrm{ang}}$, and is hence given by (cf. Eqs. (2.4)-(2.8) and Eq. (2.10)) [Zhou:2022yxp]
@@
The time-dependence of the power enters in the $\mathbf{\hat{m}}\cdot\mathbf{\hat{r}}$ term, which is periodic with a period that equals the pulsar period $P$. For simplicity in our forecast, we consider the time-averaged signal. However, we note that since the timing resolution of DSA-2000 exceeds the frequency of a typical NS with a large magnetic field (and $P\gtrsim 1$ s). We expect an improved sensitivity by weighting the signal across the rotation phase, though this requires a more detailed analysis of the signal’s time dependence, since Eq. (2) provides only an order-of-magnitude estimate on the signal amplitude. We leave such an analysis to future work.
@@
$=$ $1.4\times 10^{-5}\,\mathrm{mJy}\left(\frac{g_{a\gamma\gamma}}{10^{-12}\,\mathrm{GeV}^{-1}}\right)^{2}\left(\frac{R_{\mathrm{NS}}}{10\,\mathrm{km}}\right)^{5/2}\left(\frac{m_{a}/2\pi}{\mathrm{GHz}}\right)^{1/3}\left(\frac{B_{0}}{10^{12}\,\mathrm{G}}\right)^{5/6}\left(\frac{P}{\mathrm{sec}}\right)^{7/6}$
$\left(\frac{\rho_{\mathrm{DM}}}{0.4\,\mathrm{GeV\,cm}^{-3}}\right)\left(\frac{M_{\mathrm{NS}}}{M_{\odot}}\right)^{1/2}\left(\frac{200\,\mathrm{km/s}}{v_{0}}\right)\left(\frac{1\,\mathrm{kpc}}{d}\right)^{2}|3\cos\theta\mathbf{\hat{m}}\cdot\mathbf{\hat{r}}-\cos\theta_{m}|^{5/6}\,$, (2.12)
[…]
$\mathcal{B}\sim 10^{-5}\left(\frac{m_{a}}{2\pi}\right)=10\,\text{kHz}\left(\frac{m_{a}/2\pi}{\mathrm{GHz}}\right)\,.$ (2.13)
Here we have taken into account the interactions of axion‑sourced photons with magnetospheric plasma, which leads to a broadening of the signal bandwidth [Battye:2021xvt, Witte:2021arp, McDonald:2023shx]. The signal-to-noise ratio (SNR) is given by
@@
Figure 3: Expected radio spectral flux density from axion-photon conversions in NSs, assuming $m_{a}/(2\pi)=1$ GHz and $g_{a\gamma\gamma}=10^{-14}$ GeV-1, for each pulsar in the Australia Telescope National Facility pulsar catalog [Manchester:2004bp], magnetars in the McGill Magnetar Catalog [Olausen:2013bpa], the Galactic Center magnetar SGR J1745-2900 with $P=3.76$ s [Kennea:2013dfa, Mori:2013yda, Shannon:2013hla], and four recently identified long-period magnetars, including PSR J0901-4046 with $P=75.89$ s [Caleb:2022xyo], GLEAM-X J162759.5-523504.3 with $P=18.2$ min [Hurley-Walker:2022], GPM J1839-10 with $P=22.0$ min [Hurley-Walker:2023], and ASKAP J1935+2148, with $P=53.8$ min [Caleb:2024nrq]. Pulsars and magnetars in the catalogs are denoted by dots, while long-period magnetars are denoted by stars. The background contours in grey are derived assuming $M_{\mathrm{NS}}=1\,M_{\odot}$, $R_{\mathrm{NS}}=10$ km, $\theta=120^{\circ}$, $\theta_{m}=10^{\circ}$, $\rho_{\mathrm{DM}}=0.4$ GeV/cm3, $d=1$ kpc, and $B_{0}$ satisfying Eq. (2.5). The NSs that are not expected to produce any observable flux with these parameters are colored in black. Here we note that PSR J0901-4046 and GLEAM-X J162759.5-523504.3 are not within DSA-2000’s field of view.
@@
To detect a photon signal from axion-photon conversion, one could either target specific NSs expected to exhibit a strong axion-induced flux or measure the total flux from a population of NSs (see, e.g., Ref. [Safdi:2018oeu]). However, in the case of DSA-2000, achieving the required frequency resolution to detect the DM axion signal (cf. Eq. (2.13)) necessitates beamforming with high temporal and spectral resolution on specific sky positions. This limitation makes a collective observation of an NS population unfavorable. Therefore, for the remainder of this Section, we focus on identifying individual targets that offer the most promising sensitivity to DM axion signals.